# 3D Feature Distillation with Object-Centric Priors

### v3

@@ Abstract
Grounding natural language to the physical world is a ubiquitous topic with a wide range of applications in computer vision and robotics. Recently, 2D vision-language models such as CLIP have been widely popularized, due to their impressive capabilities for open-vocabulary grounding in 2D images. Recent works aim to elevate 2D CLIP features to 3D via feature distillation, but either learn neural fields that are scene-specific and hence lack generalization, or focus on indoor room scan data that require access to multiple camera views, which is not practical in robot manipulation scenarios. Additionally, related methods typically fuse features at pixel-level and assume that all camera views are equally informative. In this work, we show that this approach leads to sub-optimal 3D features, both in terms of grounding accuracy, as well as segmentation crispness. To alleviate this, we propose a multi-view feature fusion strategy that employs object-centric priors to eliminate uninformative views based on semantic information, and fuse features at object-level via instance segmentation masks. To distill our object-centric 3D features, we generate a large-scale synthetic multi-view dataset of cluttered tabletop scenes, spawning 15k scenes from over 3300 unique object instances, which we make publicly available. We show that our method reconstructs 3D CLIP features with improved grounding capacity and spatial consistency, while doing so from single-view RGB-D, thus departing from the assumption of multiple camera views at test time. Finally, we show that our approach can generalize to novel tabletop domains and be re-purposed for 3D instance segmentation without fine-tuning, and demonstrate its utility for language-guided robotic grasping in clutter.
@@
Keywords: Open-Vocabulary 3D Segmentation, Multi-view Feature Distillation
@@ 1 Introduction
Language grounding in 3D environments plays a crucial role in realizing intelligent systems that can interact naturally with the physical world. In the robotics field, being able to precisely segment desired objects in 3D based on open language queries (object semantics, visual attributes, affordances, etc.) can serve as a powerful proxy for enabling open-ended robot manipulation. As a result, research focus on 3D segmentation methods has seen growth in recent years [1, 2, 3, 4, 5, 6]. However, related methods fall in the closed-vocabulary regime, where only a fixed list of classes can be used as queries. Inspired by the success of open-vocabulary 2D methods  [7, 8, 9, 10], recent efforts elevate 2D representations from pretrained models  [7, 11] to 3D via distillation pipelines  [12, 13, 14, 15, 16, 17, 18, 19]. However, we identify certain limitations of existing distillation approaches. On the one hand, field-based methods  [13, 20, 16, 17, 18] offer continuous 3D feature fields, but require to be trained online in specific scenes and hence cannot generalize to novel object instances and compositions, they require a few minutes to train, and need to collect multiple camera views before training, all of which hinder their real-time applicability. On the other hand, original 3D feature distillation methods and follow up work [12, 14, 21] use room scan datasets  [22, 23] to learn point-cloud encoders, hence being applicable in novel scenes with open vocabularies. However, such approaches assume that 2D features from all views are equally informative, which is not the case in highly cluttered indoors scenes (e.g. due to partial occlusions from some view), thus leading in noisy 3D features. 2D features are also usually fused point-wise from ViT patches  [9, 10, 8] or multi-scale crops [13, 6], therefore leading to the so called “patchyness” issue [24] (see Fig. 1). The latter issue is especially impactful in robot manipulation, where precise 3D segmentation is vital for specifying robust actuation goals.
@@
To address such limitations, we revisit 2D $\rightarrow$ 3D point-based feature distillation but revise the multi-view feature fusion strategy to enhance the quality of the target 3D features. In particular, we inject both semantic and spatial object-centric priors into the fusion strategy, in three ways: (i) We obtain object-level 2D features by isolating object instances in each camera view from their 2D segmentation masks, (ii) we fuse features only at corresponding 3D object regions using 3D segmentation masks, (iii) we leverage object-level semantic information to devise an informativeness metric, which is used to weight the contribution of views and eliminate uninformative ones. Extensive ablation studies demonstrate the advantages of object-centric fusion compared to vanilla approaches. To train our method, we require a large-scale cluttered indoors dataset with many views per scene, which is currently not existent. To that end, we build MV-TOD (Multi-View Tabletop Objects Dataset), consisting of $\sim 15k$ Blender scenes from more than $3.3k$ unique 3D object models, for which we provide $73$ views per scene with $360^{\circ}$ coverage, further equipped with 2D/3D segmentations, 6-DoF grasps and textual object-level annotations. We use MV-TOD to distill our object-centric 3D CLIP  [7] features into a 3D representation, which we call DROP-CLIP (Distilled Representations with Object-centric Priors from CLIP). Our 3D encoder operates in partial point-clouds from a single RGB-D view, thus departing from the requirement of multiple camera images at test time, while offering real-time inference capabilities. We demonstrate that our learned 3D features achieve high grounding performance and segmentation crispness, while significantly outperforming previous 2D open-vocabulary approaches in the single-view setting. Further, we show that they can be leveraged zero-shot in novel tabletop domains, as well as be used out-of-the-box for 3D instance segmentation.
@@ 2 Multi-View Tabletop Objects Dataset
Figure 2: (Left:) Example generated Blender scene, multi-view image coverage and annotations included in our dataset. (Right:) Automatic semantic annotation generation with large vision-language models.
@@
Table 1: Comparisons between MV-TOD and existing 3D datasets.
Existing 3D datasets mainly focus on indoor scenes in room layouts [33, 22, 26] and related language annotations typically cover closed-set object categories (e.g. furniture) and spatial relations [1, 2, 27, 34, 28], which are not practical for robot manipulation tasks, where cluttered tabletop scenarios and open-vocabulary language are of key importance. On the other hand, recent grasp-related efforts collect cluttered tabletop scenes, but either lack language annotations [30, 35, 29] or connect cluttered scenes with language but only for 4-DoF grasps with RGB data  [31, 32]. Further, most of such datasets lack dense multi-view scene coverage, granting them non applicable for 2D $\rightarrow$ 3D feature distillation, where we require multiple images from each scene to extract 2D features with a foundation model. To cover this gap, we propose MV-TOD, a large-scale synthetic dataset with cluttered tabletop scenes featuring dense multi-view coverage and rich language annotations at the object level. We generate a total of $15k$ scenes in Blender [36], comprising of 3379 unique object models, $99$ collected by us and the rest filtered from ShapeNet-Sem model set  [37]. The dataset features $149$ object categories, each of which includes multiple instances that vary in fine-grained details. For each object instance, we leverage GPT-4-Vision [38] to generate open-set descriptions from various perspectives, including category, color, material, state, utility, affordance, etc, which spawn over $670k$ unique referring instance queries (see Fig. 2-right and Appendix A). For each scene, we provide 2D/3D segmentation masks, 6D object poses, as well as a set of semantic concepts for each appearing object instance. Additionally, we include 6-DoF grasp annotations for each object model, originating from the ACRONYM dataset  [35]. To the best of our knowledge, MV-TOD is the first dataset to combine 3D cluttered tabletop scenes with open-vocabulary language and 6-DoF grasp annotations, which we hope will accelerate future research.
3 Methodology
@@
Our goal is to distill multi-view 2D CLIP features into a 3D representation, while employing an object-centric feature fusion strategy to ensure high quality 3D features. Our overall pipeline is illustrated in Fig 3. We first introduce traditional multi-view feature fusion (Sec. 3.1), present our variant with object-centric priors (Sec. 3.2) and discuss our feature distillation method (Sec. 3.3).
3.1 Multi-view 2D Feature Fusion
We assume access to a dataset of 3D scenes, where each scene is represented through a set of $\mathcal{V}$ posed RGB-D views of size $H\times W$: $\left\{I_{v}\in\mathbb{R}^{H\times W\times 3},\,D_{v}\in\mathbb{R}^{H\times W},\,T_{v}\in\mathbb{R}^{4\times 4}\right\}_{v=1}^{\mathcal{V}}$, where $T_{v}$ the camera pose from view $v$. For each scene, we first obtain the full point-cloud $P\in\mathbb{R}^{M\times 3}$ along with a 2D-3D correspondence map $\mathcal{M}_{v}\in\mathbb{R}^{M\times 2}$, mapping each point $\mathbf{x}_{i},\,i=1,\dots,M$ to a pixel location $\mathbf{u}_{v,i}=(u_{x},u_{y})^{T}$ in image view $I_{v}$. RGB images are fed to a pretrained image model $f^{2D}:\mathbb{R}^{H\times W\times 3}\rightarrow\mathbb{R}^{H\times W\times C}$ [9, 10, 8] to obtain pixel-level 2D features of size C: $Z^{2D}_{v}=f^{2D}(I_{v})$, which are then back-projected to 3D points via:
@@ 3.1 Multi-view 2D Feature Fusion
To fuse 2D features across views $Z^{3D}\in\mathbb{R}^{M\times C}$, previous works  [12, 6, 14, 21] use average pooling: $\mathbf{z}_{i}^{3D}=\frac{1}{\mathcal{V}}\sum_{v=1}^{\mathcal{V}}\mathbf{z}_{v,i}^{2D}$ (see Appendix B for a comprehensive overview). In essence, this method assumes that all views are equally informative for each point, as long as the point is visible from that view. We suggest that naively average pooling 2D features for each point leads to sub-optimal 3D features, as noisy, uninformative views contribute equally, therefore “polluting" the overall representation. We then propose to instead use a generalized version relying on weighted average:
@@
where $\omega_{v,i}\in\mathbb{R}$ is a scalar weight that represents the informativeness of view $v$ for point $i$. In the next subsection, we describe how to use text data to dynamically compute an informativeness weight for each view based on semantic object-level semantic information. Additionally, vanilla point-wise fusion with pixel-level 2D features leads to non-crisp segmentations and fuzzy object boundaries. To resolve this, we propose to also leverage dense spatial information, i.e., instance-wise 2D/3D segmentation masks, which are used for both: (a) obtaining robust object-level 2D CLIP features from each view, and (b) fusing features only at the points corresponding to the 3D object region.
@@ 3.2 Employing Object-Centric Priors
Let $\left\{S_{v}^{2D}\in\{0,1\}^{N\times H\times W}\right\}_{v=1}^{\mathcal{V}}$ be 2D instance-wise segmentation masks for each scene, where $N$ the total number of scene objects. We aggregate the 2D masks to obtain $S^{3D}\in\{0,1\}^{M\times N}$, such that for each point $i$ we can retrieve the corresponding object instance $n_{i}=\texttt{argmax}_{\nu}\;S^{3D}_{i,\nu}$.
@@
Object-level 2D CLIP features For obtaining object-level 2D CLIP features, we isolate the pixels for each object $n$ from each view $v$ from $S^{2D}_{v,n}$ and crop a bounding box around the mask from $I_{v}$: $\textbf{z}^{2D}_{v,n}=f^{2D}_{cls}\left(\texttt{cropmask}(I_{v},\,S^{2D}_{v,n})\right)$ (see Appendix C for ablations in CLIP visual prompts). Here we use $f^{2D}_{cls}:\mathbb{R}^{h_{n}\times w_{n}\times 3}\rightarrow\mathbb{R}^{C}$, i.e., only the [CLS] feature of CLIP’s ViT encoder, to represent an object crop of size $h_{n}\times w_{n}$. We can now define our metric from equation (3) also at object-level:
@@
where each view is weighted by its semantic informativeness metric $G_{v,n}$, as well as optionally a visibility metric $\Lambda_{v,n}=\sum_{\textbf{}}S^{2D}_{v,n}$ that measures the number of pixels from $n$-th object’s mask that are visible from view $v$  [6]. We finally reconstruct the full feature-cloud $Z^{3D}\in\mathbb{R}^{M\times C}$ by equating each point’s feature to its corresponding 3D object-level one via: $\mathbf{z}^{3D}_{i}=\mathbf{z}^{3D}_{n_{i}},\;n_{i}=\texttt{argmax}_{\nu}\;S^{3D}_{i,\nu}$.

### v4

@@ Abstract
Grounding natural language to the physical world is a ubiquitous topic with a wide range of applications in computer vision and robotics. Recently, 2D vision-language models such as CLIP have been widely popularized, due to their impressive capabilities for open-vocabulary grounding in 2D images. Subsequent works aim to elevate 2D CLIP features to 3D via feature distillation, but either learn neural fields that are scene-specific and hence lack generalization, or focus on indoor room scan data that require access to multiple camera views, which is not practical in robot manipulation scenarios. Additionally, related methods typically fuse features at pixel-level and assume that all camera views are equally informative. In this work, we show that this approach leads to sub-optimal 3D features, both in terms of grounding accuracy, as well as segmentation crispness. To alleviate this, we propose a multi-view feature fusion strategy that employs object-centric priors to eliminate uninformative views based on semantic information, and fuse features at object-level via instance segmentation masks. To distill our object-centric 3D features, we generate a large-scale synthetic multi-view dataset of cluttered tabletop scenes, spawning 15k scenes from over 3300 unique object instances, which we make publicly available. We show that our method reconstructs 3D CLIP features with improved grounding capacity and spatial consistency, while doing so from single-view RGB-D, thus departing from the assumption of multiple camera views at test time. Finally, we show that our approach can generalize to novel tabletop domains and be re-purposed for 3D instance segmentation without fine-tuning, and demonstrate its utility for language-guided robotic grasping in clutter.
@@ 1 Introduction
Language grounding in 3D environments plays a crucial role in realizing intelligent systems that can interact naturally with the physical world. In the robotics field, being able to precisely segment desired objects in 3D based on open language queries (object semantics, visual attributes, affordances, etc.) can serve as a powerful proxy for enabling open-ended robot manipulation. As a result, research focus on 3D segmentation methods has seen growth in recent years [1, 2, 3, 4, 5, 6]. However, related methods fall in the closed-vocabulary regime, where only a fixed list of classes can be used as queries. Inspired by the success of open-vocabulary 2D methods  [7, 8, 9, 10], recent efforts elevate 2D representations from pretrained image models  [7, 11] to 3D via distillation pipelines  [12, 13, 14, 15, 16, 17, 18, 19]. In this work, we identify several limitations of existing distillation approaches. On the one hand, field-based methods  [13, 20, 16, 17, 18] offer continuous 3D feature fields, but require to be trained online in specific scenes and hence cannot generalize to novel object instances and compositions, they require a few minutes to train, and need to collect multiple camera views before training, all of which hinder their real-time applicability. On the other hand, original 3D feature distillation methods and follow up work [12, 14, 21] use room scan datasets  [22, 23] to distill 2D features fused from multiple views with point-cloud encoders. The distilled features maintain the open-set generalizability of the pretrained model, therefore granting such methods applicable in novel scenes with open vocabularies. However, such approaches assume that 2D features from all views are equally informative, which is not the case in natural indoors scenes, where due to partial visibility and clutter, certain views will lead to noisy representations. 2D features are also typically fused point-wise from ViT patches  [9, 10, 8] or multi-scale crops [13, 6], therefore leading to the so called “patchyness” issue [24] (see Fig. 1), where features computed in patches / crops that involve multiple objects lead to fuzzy segmentation boundaries. The latter issue is especially impactful in robot manipulation, where precise 3D segmentation is vital for specifying robust actuation goals.
@@
To address such limitations, in this work, we revisit 2D $\rightarrow$ 3D feature distillation with point-cloud encoders, but revise the multi-view feature fusion strategy to enhance the quality of the target 3D features. In particular, we inject both semantic and spatial object-centric priors into the fusion strategy, in three ways: (i) We obtain object-level 2D features by isolating object instances in each camera view from their 2D segmentation masks, (ii) we fuse features only at corresponding 3D object regions using their 3D segmentation masks, (iii) we leverage dense object-level semantic information to devise an informativeness metric, which is used to weight the contribution of views and eliminate uninformative ones. Extensive ablation studies demonstrate the advantages of our proposed object-centric fusion strategy compared to vanilla approaches. To train our method, we require a large-scale cluttered indoors dataset with dense number of views per scene, which is currently not existent. To that end, we build MV-TOD (Multi-View Tabletop Objects Dataset), consisting of $\sim 15k$ Blender scenes from more than $3.3k$ unique 3D object models, for which we provide $73$ views per scene with $360^{\circ}$ coverage, further equipped with 2D/3D segmentations, 6-DoF grasps and semantic object-level annotations. We use MV-TOD to distill the object-centric 3D CLIP  [7] features acquired via our fusion strategy into a 3D representation, which we call DROP-CLIP (Distilled Representations with Object-centric Priors from CLIP). Our 3D encoder operates in partial point-clouds from a single RGB-D view, thus departing from the requirement of multiple camera images at test time, while offering real-time inference capabilities. By imposing the same 3D features as distillation targets for a large number of diverse views, we encourage DROP-CLIP to learn a view-invariant 3D representation. We demonstrate that our learned 3D features surpass previous 3D open-vocabulary approaches in semantic and referring segmentation tasks in MV-TOD, both in terms of grounding accuracy and segmentation crispness, while significantly outperforming previous 2D approaches in the single-view setting. Further, we show that they can be leveraged zero-shot in novel tabletop datasets that contain real-world scenes with unseen objects and new vocabulary, as well as be used out-of-the-box for 3D instance segmentation tasks, performing competitively with established segmentation approaches without fine-tuning.
@@ 2 Multi-View Tabletop Objects Dataset
Figure 2: MV-TOD Overview: Example generated scene, source multi-view RGB=D images and scene annotations (left). Automatic semantic annotation generation with VLMS (right).
@@
Table 1: Comparisons between MV-TOD and existing datasets.
Existing 3D datasets mainly focus on indoor scenes in room layouts [33, 22, 26] and related annotations typically cover closed-set object categories (e.g. furniture) [1, 2, 27, 34, 28], which are not practical for robot manipulation tasks, where cluttered tabletop scenarios and open-vocabulary language are of key importance. Alternatively, recent grasp-related research efforts collect cluttered tabletop scenes, but either lack language annotations [30, 35, 29] or connect cluttered scenes with language but only for 4-DoF grasps with RGB data  [31, 32], hence lacking crucial 3D information. Further, all existing datasets lack dense multi-view scene coverage, granting them non applicable for 2D $\rightarrow$ 3D feature distillation, where we require multiple images from each scene to extract 2D features with a foundation model. To cover this gap, we propose MV-TOD, a large-scale synthetic dataset with cluttered tabletop scenes featuring dense multi-view coverage, segmentation masks, 6-DoF grasps and rich language annotations at the object level (see Fig. 2). Table 1 summarizes key differences between MV-TOD and existing grounding / grasping datasets.
MV-TOD contains approximately $15k$ scenes generated in Blender [36], comprising of 3379 unique object models, $99$ collected by us and the rest filtered from ShapeNet-Sem model set [37]. The dataset enumerates $149$ object categories featuring typical household objects (kitchenware, food, electronics etc.), each of which includes multiple instances that vary in fine-grained details such as color, texture, shape etc. For each object instance, we leverage modern vision-language models such as GPT-4-Vision [38] to generate textual annotations referring to various object attributes, including category, color, material, state, utility, brand, etc., spawning over $670k$ unique referring instance queries. We refer the reader to Appendix A.1 for details on object statistics and scene generation implementation. For each scene, we provide 73 uniformly distributed views, 2D / 3D instance segmentation masks, 6D object poses, as well as a set of referring expressions sampled from the object-level semantic annotations. Additionally, we provide collision-free 6-DoF grasp poses for each scene object, originating from the ACRONYM dataset  [35]. In this paper, we leverage the dense multi-view coverage of MV-TOD for 2D $\rightarrow$ 3D feature distillation. However, given the breadth of labels in MV-TOD, we believe it can serve as a resource for several 3D vision and robotics downstream tasks, including instance segmentation, 6D pose estimation and 6-DoF grasp synthesis. To the best of our knowledge, MV-TOD is the first dataset to combine 3D cluttered scenes with multi-view images, open-vocabulary language and 6-DoF grasp annotations.
3 Distilled Representations with Object-Centric Priors
@@
Our goal is to distill multi-view 2D CLIP features into a 3D representation, while employing an object-centric feature fusion strategy to ensure high quality 3D features. Our overall pipeline is illustrated in Fig 3. We first introduce traditional multi-view feature fusion (Sec. 3.1), present our variant with object-centric priors (Sec. 3.2), discuss feature distillation training (Sec. 3.3) and describe how to perform inference for downstream open-vocabulary 3D grounding tasks (Sec. 3.4).
3.1 Multi-view Feature Fusion
We assume access to a dataset of 3D scenes, where each scene is represented through a set of $\mathcal{V}$ posed RGB-D views $\left\{I_{v}\in\mathbb{R}^{H\times W\times 3},\,D_{v}\in\mathbb{R}^{H\times W},\,T_{v}\in\mathbb{R}^{4\times 4}\right\}_{v=1}^{\mathcal{V}}$, with $H\times W$ denoting the image resolution, $\mathcal{V}$ the total number of views, and $T_{v}$ the transformation matrix from each camera’s viewpoint $v$ with respect to a global reference frame, such as the center of the tabletop. A projection matrix $K_{v}$ representing each camera’s intrinsic parameters is also given. For each scene we reconstruct the full point-cloud $P\in\mathbb{R}^{M\times 3}$ by aggregating all depth images $D_{v}$, after projecting them to 3D with the camera intrinsics $K_{v}$ and transforming to world frame with $T_{v}^{-1}$. To remove redundant points, we voxelize the aggregated point-cloud with a fixed voxel size resolution $d^{3}$, resulting in $M$ total points. Our goal is to obtain a feature-cloud $Z^{3D}\in\mathbb{R}^{M\times C}$, where $C$ is the dimension of the representations provided by the pretrained image model, fused across all views.
2D feature extraction We pass each RGB view to a pretrained image model $f^{2D}:\mathbb{R}^{H\times W\times 3}\rightarrow\mathbb{R}^{H\times W\times C}$ to obtain pixel-level features $Z^{2D}_{v}=f^{2D}(I_{v})$. Any ViT-based vision foundation model (e.g. DINO-v2  [11]) can be chosen, but we focus on CLIP  [7], since we want our 3D representation to be co-embedded with language, as to enable open-vocabulary grounding. However, vanilla CLIP features are restrained to image-level, whereas we require dense pixel-level features to perform multi-view fusion. To obtain pixel-wise features, previous works explore fine-tuned CLIP models  [12, 15] such as OpenSeg  [9] or LSeg  [10], multi-scale crops from anchored points in the image frame  [13, 6, 21] or MaskCLIP  [8, 16], which provides patch-level text-aligned features from CLIP’s ViT encoder without additional training. All approaches are compatible with our framework (ablations in Sec. 4.1).
2D-3D correspondence Given the $i$-th point in $P$, $\textbf{x}_{i}=\left(x,y,z\right),\,i=1,\dots,M$, we first back-project to each camera view $v$ using: $\tilde{\mathbf{u}}_{v,i}\doteq\mathcal{M}_{v}(\mathbf{x}_{i})=K_{v}\cdot T_{v}\cdot\tilde{\textbf{x}}_{i}$, where $\tilde{\mathbf{u}}=\left(u_{x},u_{y},u_{z}\right)^{T}$ and $\tilde{\mathbf{x}}=\left(x,y,z,1\right)^{T}$ homogeneous coordinates in 2D camera frame and 3D world frame respectively, and $\mathbf{u}=(u_{x},u_{y})^{T}$. The 2D feature for each back-projected point $\textbf{z}^{2D}_{v,i}\in\mathbb{R}^{C}$ is then given by:
@@ 3.1 Multi-view Feature Fusion
For each view, we eliminate points that fall outside of a camera view’s FOV by considering only the pixels: $\left\{\tilde{\textbf{u}}_{v}=\left(u_{x},u_{y},u_{z}\right)^{T}\in\mathcal{M}_{v}(P)\mid u_{z}\neq 0,\;u_{x}/u_{z}\in[0,W),\;u_{y}/u_{z}\in[0,H)\right\}$. It is further important to maintain only points that are visible from each camera view, as a point might lie within the camera’s FOV but in practise be occluded by a foreground object. To eliminate such points, we follow  [12, 6] and compare the back-projected z coordinate $u_{z}$ with the sensor depth reading $D_{v}(u_{x},u_{y})$. We maintain only points that satisfy: $\left|u_{z}-D_{v}({u_{x},u_{y}})\right|\leq c_{thr}$, where $c_{thr}$ a fixed hyper-parameter. We compose the FOV and occlusion filtering to obtain a visibility map $\Lambda_{v,i}\in\{0,1\}^{\mathcal{V}\times M}$, which determines whether point $i$ is visible from view $v$.
Fusing point-wise features Obtaining a 3D feature for each point $i=1,\dots,M$ is achieved by fusing back-projected 2D features $Z^{2D}_{v}$ with weighted-average pooling:
@@
where $\omega_{v,i}\in\mathbb{R}$ a scalar weight that represents the importance of view $v$ for point $i$. In practise, previous works consider $\omega_{v,i}=\Lambda_{v,i}$ [12], a binary weight for the visibility of each point. In essence, this method assumes that all views are equally informative for each point, as long as the point is visible from that view.
We suggest that naively average pooling 2D features for each point leads to sub-optimal 3D features, as noisy, uninformative views contribute equally, therefore “polluting" the overall representation. In our work we propose to decompose $\omega_{v,i}=\Lambda_{v,i}\cdot G_{v,i}$, where $G_{v,i}\in\mathbb{R}^{\mathcal{V}\times M}$ an informativeness weight that measures the importance of each view for each point. In the next subsection, we describe how to use text data to dynamically compute an informativeness weight for each view based on semantic object-level information, as well as how to perform object-wise instead of point-wise fusion.
@@ 3.2 Employing Object-Centric Priors
Let $\left\{S_{v}^{2D}\in\{0,1\}^{N\times H\times W}\right\}_{v=1}^{\mathcal{V}}$ be view-aligned 2D instance-wise segmentation masks for each scene, where $N$ the total number of scene objects. We aggregate the 2D masks to obtain $S^{3D}\in\{1,\dots,N\}^{M}$, such that for each point $i$ we can retrieve the corresponding object instance $n_{i}=S^{3D}_{i}$.
@@
Object-level 2D CLIP features For obtaining object-level 2D CLIP features, we isolate the pixels for each object $n$ from each view $v$ from $S^{2D}_{v,n}$ and crop a bounding box around the mask from $I_{v}$: $\textbf{z}^{2D}_{v,n}=f^{2D}_{cls}\left(\texttt{cropmask}(I_{v},\,S^{2D}_{v,n})\right)$ (see Appendix A.3 for ablations in CLIP visual prompts). Here we use $f^{2D}_{cls}:\mathbb{R}^{h_{n}\times w_{n}\times 3}\rightarrow\mathbb{R}^{C}$, i.e., only the [CLS] feature of CLIP’s ViT encoder, to represent an object crop of size $h_{n}\times w_{n}$. We can now define our metric from equation (3) also at object-level:
@@
where each view is weighted by its semantic informativeness metric $G_{v,n}$, as well as optionally a visibility metric $\Lambda_{v,n}=\sum_{\textbf{}}S^{2D}_{v,n}$ that measures the number of pixels from $n$-th object’s mask that are visible from view $v$  [6]. We finally reconstruct the full feature-cloud $Z^{3D}\in\mathbb{R}^{M\times C}$ by equating each point’s feature to its corresponding 3D object-level one via: $\mathbf{z}^{3D}_{i}=\mathbf{z}^{3D}_{n_{i}},\;n_{i}=S^{3D}_{i}$.